OpenStudy (anonymous):

The perimeter of a rectangle is 46 centimeters. Its width is 5 centimeters. Find the length. a. 19 cm c. 16 cm b. 18 cm d. 15 cm

OpenStudy (texaschic101):

formula : P = 2L + 2W now sub your information in the formula 46 = 2L + 2(5) 46 = 2L + 10 46 - 10 = 2L 36 = 2L 18 = L check... P = 2L + 2W 46 = 2(18) + 2(5) 46 = 36 + 10 46 = 46 (correct) ANSWER : B. 18 cm

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

no i get it now thnx

OpenStudy (texaschic101):

no problem :)